OpenStudy (anonymous):

if you have let's say y*(((x+5)-2)/y) can i say that this equals to (x+5)-2??

OpenStudy (cwrw238):

if the capital Y = y then yes you can

OpenStudy (anonymous):

\[y*\frac{ x+5 }{ y }-2 =(x+5)-2\]

OpenStudy (anonymous):

not sure what the capital Y means or what's the difference

OpenStudy (aum):

The y's cancel out (assuming \(y \ne 0\). So what you have is correct.

OpenStudy (anonymous):

now what if the original equation is y*c and i'm substituting \[\frac{ x+5 }{ y }-2 \] for c

OpenStudy (aum):

\[ y * \left ( \frac{ x+5 }{ y }-2\right) = x + 5 - 2y \]